In a survey of 385 students at Broward College, North Campus, results revealed that 154 students...

  1. In a survey of 385 students at Broward College, North Campus, results revealed that 154 students were born outside of the USA. Construct a 96% confidence interval estimate for the proportion of students who were born outside the USA. Show your calculations on a separate sheet of paper for p-hat. Q-hat, the margin of error and the construction of the confidence interval. Also, give your interpretation of the confidence interval estimate. SHOW WORK
Answer #1

Sample proportion = 154 / 385 = 0.40

= 1 - = 1 - 0.4 = 0.6

Margin of error = Z/2 * sqrt( * / n)

= 2.0537 * sqrt( 0.4 * 0.6 / 385)

= 0.0513

96% Confidence interva for p is

- ME < p < + ME

0.4 - 0.0513 < p < 0.4 + 0.0513

0.379 < p < 0.451

96% CI is ( 0.379 , 0.451)

Interpretation - we are 96% confident that the proportion of students who born outside the USA is

between 0.379 and 0.451.